OpenStudy (anonymous):

solve the differential equation dy/dx=(x-2)2

OpenStudy (jhannybean):

\[\large \frac{dy}{dx}=(x-2)^2\] this is a separable equation, sowe can transfer the dx over to the right side to integrate

OpenStudy (jhannybean):

\[\large \int\limits{dy}=\int\limits (x-2)^2dx\] are you able to integrate this?

OpenStudy (anonymous):

yeah thank you so much

OpenStudy (jhannybean):

awesome :)